### v2

# Image segmentation of treated and untreated tumor spheroids by Fully Convolutional Networks

Abstract
Background and objectives
Multicellular tumor spheroids (MCTS) are advanced cell culture systems for assessing the impact of combinatorial radio(chemo)therapy. They exhibit therapeutically relevant in-vivo-like characteristics from three-dimensional (3D) cell-cell and cell-matrix interactions to radial pathophysiological gradients related to proliferative activity and nutrient/oxygen supply, altering cellular radioresponse. The therapeutic outcome of MCTS is most frequently monitored over time via brightfield microscopy imaging. Classical yet non-curative analytical readouts relate to spheroid size and volume growth kinetics. State-of-the-art assays also quantify long-term curative endpoints based on collected image time series from large treated spheroid populations per irradiation dose and treatment arm. Here, spheroid control probabilities are documented analogous to in-vivo tumor control probabilities based on Kaplan-Meier curves. These analyses require spheroid segmentation to extract relevant structural information from the images, e.g., diameter, area, volume and circularity. This laborious process entails manual inspection and analysis of up to 100.000 images per treatment arm. While several image analysis algorithms are available for spheroid segmentation, they all focus on compact MCTS with clearly distinguishable outer rim throughout growth. However, treated MCTS may partly be detached and destroyed and are usually obscured by dead cell debris. Thus, this work aims to develop an advanced automatic segmentation that also covers the previously neglected cases of spheroids obscured by debris.
Methods
We successfully train two Fully Convolutional Networks, U-Net and HRNet, and optimize their hyperparameters to develop an automatic segmentation for both untreated and treated MCTS. We systematically validate the automatic segmentation on larger, independent data sets of spheroids derived from two human head-and-neck cancer cell lines.
Results
We find an excellent overlap between manual and automatic segmentation for most images, quantified by Jaccard indices at around 90%. For images with smaller overlap of the segmentations, we demonstrate that this error is comparable to the variations across segmentations from different biological experts (inter-observer variability), suggesting that these images represent biologically unclear or ambiguous cases.
Conclusion
The developed automatic segmentation cannot only be used directly but also integrated into existing spheroid analysis pipelines and tools. This facilitates the analysis of 3D spheroid assay experiments and contributes to the reproducibility and standardization of this preclinical in-vitro model.
I Introduction
One of the most challenging problems in oncology is the development of therapeutic approaches for tumor growth suppression and eradication, as well as designing and optimizing treatment protocols. In this context, three-dimensional (3D) multicellular tumor spheroids (MCTS) are an advocated pre-clinical, in-vitro model to systematically study possible means of tumor suppression, assess the curative effect of combinatorial radio(chemo)therapy, and predict the response of in-vivo tumors FraMicAlaKunVosLan2023 ; BruZieRivOelHaa2019 ; BruRivBoxOelHaa2020 ; RifHeg2017 ; LeeGriHarMcI2016 ; HirMenDitWesMueKun2010 ; FriEbnKun2007 ; KunFreHofEbn2004 . In contrast to two-dimensional (2D) clonogenic survival assays, which are known to reflect the therapeutic response of cancer cells in the tissue comparatively poorly Peretal2020 , multicellular spheroids are reproducible 3D avascular clusters of several thousand tumor cells without or in advanced settings with stromal cell compartments mimicking the pathophysiological milieu of tumor microareas or micrometastases. Due to their more or less radial 3D histomorphology and structure, they exhibit many characteristic features affecting tumor growth dynamics, including 3D reciprocal cell-cell and cell-extracellular-matrix interactions as well as metabolic gradients of oxygen, nutrients and waste products, which strongly impair the cells’ proliferative activity and therapy response RifHeg2017 ; LeeGriHarMcI2016 ; HirMenDitWesMueKun2010 ; FriEbnKun2007 ; KunFreHofEbn2004 , in particular oxygen deficiency (hypoxia), which is associated with substantial radioresistance HowAlp1957 ; MuzPueAzaAza2015 ; RakEscFurMeiLar2021 ; LiWisOkoEtal2022 . After a long phase in which spheroids were used only in specialized laboratories, methodological advances in serial culturing and live imaging have led to an exponential spread of this in-vitro model system over the past two decades KunFreHofEbn2004 ; FriEbnKun2007 ; Peietal2021 .
While 3D tumor spheroids provide a physiologically more realistic in-vitro framework to study tumor growth dynamics and therapeutic outcomes, the analysis of the experiments is much more complex than for traditional 2D cultures. 3D tumor spheroids’ dynamics with and without treatment are most frequently monitored via microscopy imaging. State-of-the-art long-term spheroid-based assays assess therapy response from these image time series by classifying each spheroid within a population as either controlled or relapsed based on their growth kinetics. By averaging the therapeutic response over ensembles of spheroids for each treatment dose, the spheroid control probabilities and spheroid control doses are computed as analytical endpoints CheManLehSorLoeYuDubBauKesStaKun2021 ; CheMicEckWonMenKraMcLKun2021 ; HinIngKaeLoeTemKoeDeuVovStaKun2018 ; VynBobGarDitStaKun2012 , analogous to the tumor control dose employed in in-vivo radiotherapy experiments with tumor-bearing mice KumLoeGurHerYarEicBauKraJen2021 ; BauPetKra2005 . Moreover, time points of relapse are identified to quantify growth recovery in terms of Kaplan-Meier curves. These metrics depend on the spheroid type (cell line), the size of the spheroids at the onset of treatment, and the applied treatment and often require the extraction of growth curves, i.e., spheroid volume over time, from the image time series of all individual spheroids.
The associated data analysis poses a significant, interdisciplinary challenge: A single, typical experimental series of a long-term spheroid-based assay, i.e., 2-3 cell models at ten different doses with 3-5 agents and 30 spheroids per treatment arm monitored up to 60 days after treatment, generates up to $10^{5}$ images. In the past, several tools were developed for spheroid image analysis HoqWinLovAve2013 ; IvaParWalAleAshGelGar2014 ; MonFerDucArgDegYou2016 ; RueSchHinDeZWalAreEli2017 ; MorPalSanTorPapPilSpiGum2017 ; BouVosTitCruMicFer2018 , including SpheroidSizer Cheetal2014 , AnaSP Pic2015 ; PicPeiStePyuTumTazWevTesMarCas2023 , TASI HouKonBraMarCoo2018 , SpheroidJ Lacetal2021 , and INSIDIA MorPalSanTorPapPilSpiGum2017 ; PerRosFriPieBarParCiaMorPapSpiPal2022 . The most crucial part of spheroid image analysis is the identification of the set of pixels in an image that corresponds to the spheroid, as this forms the basis for the extraction of spheroid characteristics like diameter, volume, and circularity as well as other morphologic features. This identification or classification of pixels in a given image is denoted semantic image segmentation Jad2020 and represents the greatest challenge in spheroid image analysis. Previous spheroid segmentations were based on classical segmentation techniques, including thresholding methods (Yen YenChaCha1995 , Otsu Ots1979 ; Pic2015 ), watershed algorithm RoeMei2000 , shape-based detections (Hough transform algorithms DudHar1972 , active contour models/snake  CasKimSap1997 ), or edge detection (Canny, Sobel Lacetal2021 ), and are also available as ImageJ plugins IvaParWalAleAshGelGar2014 ; MorPalSanTorPapPilSpiGum2017 , Matlab packages HouKonBraMarCoo2018 ; Cheetal2014 ; Pic2015 , or dedicated segmentation programs CisOanStuReg2016 ; VinGowBoxPatZimCouLomMenHarEcc2012 ; CorCikGus2012 . Since the characteristics of spheroid images, including the size, shape, and texture of the actual spheroid, vary with cell line, treatment, and microscopy method (e.g., brightfield SadKarWae2017 ; Lacetal2021 ; MorPalSanTorPapPilSpiGum2017 ; PerRosFriPieBarParCiaMorPapSpiPal2022 ; PicPeiStePyuTumTazWevTesMarCas2023 , fluorescence Lacetal2021 ; CelRizBlaMasGliPogHas2014 ; PicTesZanBev2017 ; MorPalSanTorPapPilSpiGum2017 ; BouVosTitCruMicFer2018 ; PerRosFriPieBarParCiaMorPapSpiPal2022 , differential, interference contrast NgoYanMaoNguNgKuoTsaSawTu2023 ), most of these approaches and tools are specialized for specific experimental conditions or even selected microscopes CelRizBlaMasGliPogHas2014 ; PicTesZanBev2017 and often fail to generalize to arbitrary conditions. In recent years, data-based methods, especially deep-learning models, have been increasingly employed for spheroid segmentation to tackle this issue of generalization SadKarWae2017 ; Zaoetal2021 ; Lacetal2021 ; ZhuCheWanTanHeQinYanJinYuJinLiKet2022 ; NgoYanMaoNguNgKuoTsaSawTu2023 ; AksKatAbeSheBesBurBigAdaMonGhe2023 ; PicPeiStePyuTumTazWevTesMarCas2023 ; GarDomHerMatPas2024 . Deep-learning segmentation also allows more complex analysis, e.g., identification of time-lapse migratory patterns NgoYanMaoNguNgKuoTsaSawTu2023 or distinction of the spheroids’ core and invasive edge PerRosFriPieBarParCiaMorPapSpiPal2022 .
However, these segmentation and analysis tools have primarily been developed for experimental conditions where the resulting images are relatively clean with well-visible, unobscured MCTS. In contrast, radiotherapy, one of the most common cancer treatments, regularly causes spheroids to shed dead cells or even to detach completely. Consequently, dead cell debris often obscures the remaining shrunk MCTS and the 3D cultures putatively regrowing from surviving viable cells after detachment, see Fig. 1 for two examples. This debris can cover a much larger domain than the actual spheroid and can be locally even thicker and thus darker, which makes segmentation difficult even for human experts. Both classical segmentation techniques and previous deep-learning models typically fail in these cases and require manual adjustments. We illustrate this challenge of spheroid segmentation by applying the four most recently proposed deep-learning models Lacetal2021 ; AksKatAbeSheBesBurBigAdaMonGhe2023 ; PicPeiStePyuTumTazWevTesMarCas2023 to images with and without severe cell debris, see App. Fig. A.1 for representative examples and A.2 for statistical results.
We train deep-learning models to segment spheroid images after radiotherapy based on annotated images from a previous experiment assessing the radioresponse of human head-and-neck squamous cell carcinoma (HNSCC). After systematic optimization of hyperparameters and pre-processing for two selected network architectures, U-Net and HRNet, the automatic segmentation exhibits high accuracy for images of both treated and untreated spheroids. We further validate the automatic segmentation with the optimized U-Net on larger, independent data sets of two cell types of head-and-neck cancer exposed to both radiotherapy and hyperthermia. For the majority of images, we find excellent overlap between manual and automatic segmentation, even in the case of small, heavily obscured spheroids, see Fig. 1. In particular, we compare segmentations of different biological experts and find that imprecisions of the automatic segmentation are comparable to variations across different humans. The optimized automatic segmentation is provided in a minimal tool with a graphical user interface.
(a) Treated - 10 Gy
(b) Zoom of white box in (a)
(c) Segmentation
(d) Treated - 10 Gy
(e) Zoom of white box in (d)
(f) Segmentation
Figure 1: Representative examples of images for automatic segmentation with the optimized U-Net (blue) compared to the manual segmentation (green) for 3D tumor spheroids after treatment. The overlap with the manual segmentation is excellent for standard size and larger spheroids obscured by cell debris (top row) and sufficient for small, heavily obscured spheroids (bottom row). Displayed are (a,d) the original images, which are also the input for the U-Net, (b,d) magnified image details around the spheroids as indicated by the white box in (a,d) for visibility, and (c,e) corresponding contours from the segmentations. The metrics for evaluation of the automatic segmentation are: the top row - JCD $=7.6\%$, RDD $=3.7\%$ and RCD $=1.6\%$; bottom row - JCD$=22.7\%$, RDD$=7.5\%$ and RCD$=14.7\%$, see text for details.
II Materials
The original data is compiled from several data sets of time images of FaDu cell line spheroids, one of the two head-and-neck cancer cell types from the previously published study on combinatorial radioresponse in human head-and-neck squamous cell carcinoma spheroid models CheMicEckWonMenKraMcLKun2021 : Spheroid populations were irradiated with X-rays at single doses of either 2.5 Gy, 5.0 Gy, 7.5 Gy or 10.0 Gy, with each dose being represented approximately equally in the data set. Spheroids were imaged for up to 60 days after treatment with brightfield microscopy, i.e., at different times after treatment and with potentially different lighting conditions. 1095 spheroids were manually segmented by a biological expert (human H1) to train and test the fully convolutional networks (FCNs). We use the train-validation-test-split technique for validation during training, which is appropriate for the large datasets and commonly used to train spheroid segmentation models Lacetal2021 ; AksKatAbeSheBesBurBigAdaMonGhe2023 ; PicPeiStePyuTumTazWevTesMarCas2023 . We split the annotated data into 3 separate data sets: 883 images for training, 108 for validation, and 104 for testing (Hold-out test set). Images from a single spheroid are exclusively assigned to one of the groups to rule out unwanted correlations and data leakage between training, validation, and testing data sets. For further testing, another set of 6574 images of head-and-neck cancer (FaDu and SAS) spheroids from the same study CheMicEckWonMenKraMcLKun2021 , treated with different combinations and doses of X-ray irradiation and hyperthermia, is manually segmented by several independent biological experts (humans H2-H5). Note that while the automatic segmentation is performed on individual images, biological experts always take precedent and subsequent images of a spheroid into account for manual segmentation.
The original images were taken with a Zeiss Axiovert 200M at a resolution of 1300 × 1030 pixels, representing $2.04\ \mu$m/pixel with 16-bit gray-scale per pixel CheMicEckWonMenKraMcLKun2021 . Images are converted to 8-bit images to ensure compatibility with the employed libraries FastAi HowGug2020 and SemTorch Cas2020 . While this conversion, in principle, reduces the contrast level, the effect is negligible as only a small fraction of the 16-bit range is utilized during imaging, e.g., the mean gray-scale value in the images is $1300\pm 130$, and values are rescaled such that smallest gray-scale value of the original image becomes 0 and the biggest value becomes 255.
(a) Valid spheroid segmentation
(b) Ambiguous spheroid segmentation
(c) Invalid spheroid segmentation
Figure 2: Exemplary evaluation of the automatic segmentation (blue) with respect to the manual one (green) for (a) standard case and (b),(c) rare artifacts, see text for details: (a) Correctly segmented spheroid: no contribution to ISF or ASF; JCD, RDD and RCD are well defined. (b) Excessive spheroids are detected beyond the actual spheroid: no contribution to ISF, one count added to ASF; JCD, RDD and RCD are well defined and computed for the larger, upper spheroid. (c) No overlap between automatic and manual segmentation: one count added to ISF, no contribution to ASF; JCD, RDD and RCD are set to one.
As deep-learning models, we use FCNs, a particular group of convolutional neural networks intended for semantic image segmentation LonSheDar2015 , in particular the U-Net RonFisBro2015 and the HRNet Wanetal2021 . The U-Net is one of the most established FCN frameworks. The HRNet has a special architecture compared to other typical FCN models and has already been successfully employed in the case of clean images with clearly visible spheroids Lacetal2021 . The pipeline to train the models employs FastAi HowGug2020 for the U-Net and its backbones and SemTorch Cas2020 for the HRNet. Before each training, a suitable learning rate is picked using LR Range Smi2018 . The training follows the 1-cycle policy: the learning rate is cycled once between two bounds, allowing fast convergence Smi2018 . During training we use validation-based early stopping, where the JCD is the underlying metric. The models are trained on a GPU NVIDIA GeForce RTX 3080 with a mini-batch size of 2. For training on images with higher resolution or to compare some specific backbones, online learning is used instead of mini-batches to avoid memory overflow, i.e., the batch size, the amount of data used in each sub-epoch weight change, is reduced to one.
III Methods
The segmentation is evaluated via several metrics (Section III.1: Jaccard Distance (JCD), Relative Diameter Deviation (RDD), Relative Circularity Deviation (RCD), Ambiguous Spheroid Fraction (ASF), Invalid Spheroid Fraction (ISF)) additional to the standard accuracy quantified by the Jaccard coefficient of the whole data set, which is used during training. These metrics are computed for individual images to report not only their mean values but also their deviations, and thus reliability, across the test data set. These metrics are used to optimize the hyperparameters, Section III.2, and data augmentation, Section III.3, of the selected FCN models U-Net and HRNet for maximal accuracy.
III.1 Evaluation metrics
Several metrics are employed to asses the accuracy of the segmentation. The most important one is the Jaccard Distance (JCD), which measures the relative difference between two sets of pixel $P$ and $T$,
JCD$=1-\text{IoU}=1-\frac{|P\cap T|}{|P\cup T|}.$ (1)
[…]
(a) Probability heatmap
(b) Segmentation map
(c) Final polygonal chain
Figure 3: Post-processing steps for the output of the FCN model to transform probability heatmap to spheroid contour: (a) Probability heatmap as output of the FCN model. Each pixel takes probability values between 0 (black) and 1 (white) predicting the target (spheroid). Note that due to the steep gradient the gradual change from black to white is hard to see. (b) By using a threshold of 0.5, the pixels are classified into outside spheroid (black) and inside spheroid (white). (c) Contour of the spheroid border extracted as a polygonal chain (blue line displayed on original image).
It is possible that the spheroid is correctly segmented, but additionally another non-existent spheroid is detected, see Fig. 2b, where additionally to the manually segmented spheroid in the top right another contour in the bottom left is suggested by the automatic segmentation. In this case, the JCD would be below 1 but differ from 0. The fraction of these cases among all images are additionally denoted Ambiguous Spheroid Fraction (ASF). Furthermore, a JCD of 1 can mean that the spheroid is found in the wrong place or the spheroid is not found at all, see Fig. 2c. The fraction of these cases is denoted Invalid Spheroid Fraction (ISF).
In practice, only the average diameter $d_{T}=2\sqrt{|T|/\pi}$ of the segmented spheroid is extracted to estimate the three-dimensional volume $\pi d_{T}^{3}/6$ of the spheroid under the assumption of a spherical shape. Additionally, the circularity $4\pi|T|/L^{2}$ of the spheroid is computed from its area $|T|$ and perimeter $L$ to assess the validity of this assumption. Note that this is just a common assumption in the field, although even a circular projection does not imply a spherical shape. Thus, we also assess the automatic segmentation by the resulting error of these two values and measure the Relative Diameter Deviation (RDD) and Relative Circularity Deviation (RCD) between the segmentations
RDDRCD$=\frac{|c_{P}-c_{T}|}{c_{T}}$ (2)
where $c_{P}$ is either mean diameter or circularity based on the automatic segmentation and $c_{T}$ is the corresponding characteristic calculated based on the manual segmentation. In the rare case that several spheroids are detected in an image, the largest one is chosen for computation of the JCD. In the case of JCD$=1$, which corresponds to an invalid spheroid, RDD and RCD are also set to 1. The combination of the JCD, ASF, ISF, RDD, and RCD quantify the accuracy of the segmentation with the FCN, where always a value closer to zero means higher agreement between the segmentations.
The spheroid’s pixel set and perimeter are required to calculate the metrics. Thus, probability heatmaps, created by the Softmax function in the last layer of the FCN, are transformed into binary images and polygonal chains as illustrated in Fig. 3. First, the binary image is created. For this purpose, every pixel with a probability value higher or equal to 50% will be counted as a spheroid pixel and set to 1. Values of the other pixels are set to 0. The coordinates from the outer border of the emerged shape in the binary image can be extracted by the algorithm proposed in SuzBe1985 , thereby creating the final polygonal chain.
III.2 Hyperparameter optimization
Several hyperparameters and pre-processing methods can be adjusted to improve the accuracy of the selected FCN models. We successively optimize the most relevant parameters, starting from the ones with the most significant expected impact on accuracy to the one with the lowest impact, i.e., backbone, transfer learning, data augmentation, input resolution, loss function, and optimizer: The backbone defines the exact architecture of the FCN model and is accordingly the most important hyperparameter. For transfer learning, the ImageNet Imagenet data set is utilized. The data augmentation techniques are described in Section III.3. Furthermore, the impact of changing the input resolution is investigated. Afterward, the loss function for quantifying the FCN’s error is optimized, particularly by comparing distribution-based loss and a region-based loss. We pick the Dice loss for region-based loss and the Cross-Entropy loss as a distribution-based loss function. We also test the Focal loss, which is a variation of Cross-Entropy and suitable for imbalanced class scenarios Jad2020 . Finally, two different optimizers for minimizing the loss function are evaluated. One is Adam KinBa2017 , a very common optimizer, and the other one is a combination of some more modern optimizers, RAdam Liuetal2021 and Lookahead Zhaetal2019 . For the backbone, the optimization is initiated with a default setting, i.e., transfer learning, no data augmentation, input resolution half of original image, Cross-Entropy loss function, and Adam optimizer. These initial parameters are then individually optimized in consecutive order.
III.3 Data Augmentation
Data augmentation is a method to generate additional training data via transformations of the original data set. Since this increases the total amount of training data, data augmentation can improve performance and avoid overfitting Busetal2020 . In our case, every original training image is transformed once, doubling the size of the original training data set. For each image in the training data set, one of three transformations is picked randomly using Albumentations Busetal2020 , i.e., vertical flip, horizontal flip, or rotation by $180\degree$. These transformations are valid, since the spheroids do not have a specific orientation within an image. Note that each of these transformation conserves the original rectangular resolution of the image. This avoids extrapolating pixels on the border of the image, which would be necessary for arbitrary rotation angles and which can lead to additional image artifacts the FCN model would have to be trained for.
IV Results
IV.1 Hyperparameter optimization
For each hyperparameter set and pre-processing method a model is trained and the performance of these trained models is compared for the validation data set based on the proposed evaluation metrics from Section III.1. For JCD, RDD, and RCD not only the average value but also the standard deviation over the whole validation data set is reported to estimate reliability of the segmentation. For the ISF and the ASF the standard error of the mean of the corresponding binomial distribution is reported. All results of the optimization are listed in the SI (SI Figs. A.5-A.14) and briefly discussed in the following.
IV.1.1 Backbones
For the U-Net, the ResNet 34 and the VGG 19-bn architecture perform best, see App. Fig. A.5, with a JCD three times smaller than the worst performing backbones, like AlexNet or SqueezeNet1.0. The two ResNet backbones stand out with a very low ASF. Note that the accuracy of the VGG architectures is improving with increasing number of used layers. The ResNet 34 is the optimal backbone as it not only performs best in virtually every metric, but also requires less computational time than the VGG 19-bn. Note that a previous study on deep-learning models for spheroid segmentation observed a higher performance for ResNet 18 compared to VGG 19 and ResNet 50 PicPeiStePyuTumTazWevTesMarCas2023 . For the HRNet, the backbones differ in the number of kernels used and the accuracy increases with the number of kernels, see App. Fig. A.6. The W48 backbone is preferred, as for the majority of metrics it performs best and has the lowest standard deviations.
IV.1.2 Data augmentation and transfer learning
The combination of data augmentation and transfer learning leads to the best performance for both U-Net and HRNet, see App. Fig. A.7 and App. Fig. A.8. While this is in principle expected, the improvement is substantial as the JCD is more than halved when both transfer learning and data augmentation is introduced. Note that for the U-Net, transfer learning improves accuracy more than data augmentation, while the opposite is true for the HRNet.
IV.1.3 Resolution
Using input images with half the resolution $650\times 515$ of the original images leads to the best performance, see App. Fig. A.9 and App. Fig. A.10. This input resolution is already used to optimize the previous parameters. Higher resolution (3/4, 1) reduces accuracy, presumably because the increase in details makes generalization of the appearance of spheroids more difficult, and the receptive fields of the kernels are getting smaller relative to the whole image. Lower resolution (1/4) also reduces accuracy, presumably due to the loss of relevant information. Half resolution displays the best accuracy and lowest standard deviations across the validation data set.
IV.1.4 Loss
We find the Dice loss to be the optimal loss function for the U-Net, see App. Fig. A.11. Note that Dice is considered the default loss parameter when using the U-Net for cell segmentation, i.e., binary classification concerned with the accuracy of the edge. However, the advantage is much smaller than for previous hyperparameters. It does not suggest a general advantage of region-based loss for the segmentation of spheroids by the U-Net. For the HRNet, there is a more significant difference between the region-based loss and the distribution-based loss, see App. Fig. A.12. The Focal loss and the Cross-Entropy loss perform better than the Dice loss - especially, the standard deviations differ by a factor of up to three. One reason for this behavior may be that the Dice loss provides less detailed feedback during training than the distribution-based losses. Cross-Entropy loss and Focal loss achieve nearly the same results, and we choose the default Cross-Entropy loss.
IV.1.5 Optimizer
In case of the U-Net the difference between the tested optimizers is marginal, see App. Fig. A.13. We choose the slighly better performing RAdam combined with Lookahead over the default optimizer Adam. In contrast, the accuracy of the HRNet is much worse for the combined optimizer, see App. Fig. A.14. The Adam optimizer is optimal with evaluation metrics 1.5 times better than for the RAdam combined with Lookahead.
Table 1: (a) Evaluation of the segmentation with the optimized U-Net and HRNet models on the test data set shows higher accuracy of the U-Net. Note that all metrics can range between zero and one, where lower values mean higher accuracy and zero means perfect agreement with the manual segmentation. (b) Evaluation of the optimized segmentations on images of untreated (not irradiated) spheroids shows high accuracy also for the case of clean images without cell debris. The optimized U-Net and HRNet achieve almost equal accuracy. (c) The optimal hyperparameter configurations.
Model JCD RDD RCD ISF ASF
U-Net 0.062 0.020 0.040 0.000 0.010
$\pm$ 0.060 0.026 0.034 0.000 0.010
HRNet 0.072 0.028 0.036 0.000 0.010
$\pm$ 0.100 0.080 0.029 0.000 0.010
(a) Segmentation test data set
Model JCD RDD RCD ISF ASF
U-Net 0.028 0.008 0.034 0.000 0.000
$\pm$ 0.010 0.008 0.032
HRNet 0.029 0.008 0.033 0.000 0.000
$\pm$ 0.011 0.007 0.026
(b) Segmentation images without cell debris
Parameter U-Net HRNet
Backbone ResNet 34 W48
Transfer learning Yes
Data augmentation Yes
Resize factor 1/2
Loss function Dice Cross-Entropy
Optimizer RAdam & Lookahead Adam
(c) Optimized parameters
IV.2 U-Net vs HRNet
The automatic segmentations with the final optimized U-Net and HRNet model are applied to the test data set. While this test data set is not taken into account during training of the models, the segmentations display virtually the same high accuracy on these new images as for the validation data set, see see Tab. 1a and App. Fig. A.15 cf. App. Fig. A.13 and App. Fig. A.14. Mostly, the standard deviations of the metrics are higher in the test data set, though the JCD remains even with this deviation well below $20\%$. As for the validation data set, the U-Net performs slightly better than the HRNet on the test data. The JCD and RDD for the HRNet are 0.01 higher, and their standard deviation is 0.04 higher. Two representative examples of the automatic segmentation by the U-Net are shown in Fig. 1: The overlap with the manual segmentation is excellent for standard size and larger spheroids obscured by cell debris, see Fig. 1 top row, and sufficient for small, heavily obscured spheroids, see bottom row. In the latter case, the model does not segment the spheroid boundary accurately but detects the spheroid at the correct position JCD $=23\%$ with a comparable size RDD$=7.5\%$ within the much larger and often darker cloud of surrounding cell debris. In such difficult cases, similar variations are observed upon manual segmentation by different biological experts.
While the models are mainly trained on images with treated spheroids which are obscured by cell debris, they also accurately segment untreated spheroids with clear boundary and clean background, see Tab. 1b: On a test set of 100 images of untreated spheroids, the accuracy of both models is nearly the same with a JCD below $3\%$. In comparison, a recent deep-learning approach for spheroids with clear boundary and clean background with focus on generalizability to different experimental conditions and microscopes, achieved a mean JCD of $8\%$ with a standard deviation of $12\%$ for bright-field microscopic images Lacetal2021 . This suggests that the standard case of clean images without cell debris is included in our training of images with heavily obscured spheroids.
The final U-Net model has a size of 158 MB and the final HRNet model a size of 251 MB. For both models, it takes about 1.8 seconds to segment one image on the CPU (Intel(R) Core(TM) i7-4770). When the segmentation is performed serially on the GPU (NVIDIA GeForce RTX 3080) the U-Net needs only 0.03 seconds per image and the HRNet 0.08 seconds. As computation time for both models is comparable, the U-Net is chosen for the automatic segmentation due to its slightly higher accuracy in out setting. Note that in a recently published deep-learning approach, the comparison of HRNet and U-Net implied that HRNet achieved the highest accuracy Lacetal2021 .
Figure 4: Validation of the automatic segmentation with the optimized U-Net on larger, independent data sets shows high accuracy for the majority of cases. (a) JCD and (b) average radial error $\Delta r$ over diameter of the manually segmented (target) spheroid $d_{T}$ for 6574 images of FaDu (blue triangles) and SAS (green stars) spheroids treated with different combinations and doses of X-ray irradiation and hyperthermia CheMicEckWonMenKraMcLKun2021 . Manual segmentation is performed by a second biological expert (human H2, blue triangles and green stars) independently from the manual segmentation (human H1, red dots) of the training, validation and test data sets. (Results for $104$ images of test data set is displayed as red dots for comparison.) Note that the segmentation is developed only based on images of FaDu spheroids. The majority of deviations are small (JCD$<0.2$, $\Delta r<20\ \mu$m, red horizontal lines as guide to the eye), average (median) values are JCD$=0.17(0.09)\pm 0.2$, $\Delta r=25(15)\pm 32\ \mu$m for the whole data set and JCD$=0.1(0.07)\pm 0.09$, $\Delta r=18(13)\pm 17\ \mu$m for spheroids larger $d_{T}\geq 400\ \mu$m (red vertical lines as guide to the eye) than the initial, standard size of spheroids. Larger imprecisions for smaller spheroids are due to biologically difficult, unclear, or ambiguous cases, see text.
IV.3 Further independent validation
[…]
In addition to the standard check with the Hold-out test data set, we systematically validate the automatic segmentation with the optimized U-Net on larger, independent Hold-out test data sets of two cell types of head-and-neck cancer, see Fig. 4. This data set contains 6574 images of FaDu or SAS spheroids treated with different combinations and doses of X-ray irradiation and hyperthermia CheMicEckWonMenKraMcLKun2021 . These images were manually segmented by another biological expert (human H2) independent from the original training/validation/test-data set used to develop the automatic segmentation (human H1). For the majority of images, we find excellent overlap between this manual and the automatic segmentation, quantified by a JCD around 0.1, see Fig. 4a. Larger deviations are mostly observed for images with smaller spheroids, i.e., below the size of standard spheroids for treatment (diameter $370-400\ \mu$m according to CheManLehSorLoeYuDubBauKesStaKun2021 ; CheMicEckWonMenKraMcLKun2021 ; HinIngKaeLoeTemKoeDeuVovStaKun2018 ). To make the evaluation intuitive for the biological experts, who are less familiar with the JCD defined in Eq. (1), we additionally introduce a measure of the average radial error $\Delta r$ based on the automatically segmented (predicted) domain $P$ and the manually segmented (target) domain $T$
$\Delta r=\sqrt{\pi^{-1}\left(|P\cup T|-|P\cap T|\right)+\frac{d_{T}^{2}}{4}}-\frac{d_{T}}{2}$ (3)
with the average (target) diameter $d_{T}$ of the manually segmented domain. The error $\Delta r$ quantifies the thickness a circular layer with the size of the mismatched area $|P\cup T|-|P\cap T|=|P\setminus T|+|T\setminus P|$ around a circle with the target area $|T|$ would have, see App. Fig. A.16 for an illustration. Since additional $P\setminus T$ and missing areas $T\setminus P$ do not compensate each other in Eq. (3), the error $\Delta r$ is considerably larger than the radial error implied by the relative difference between target and predicted diameter RDD, see App. Fig. A.17. The average radial error $\Delta r$ reported in Fig. 4b assesses the segmentation analogous to the JCD with majority of errors $\Delta r<20\ \mu$m and larger deviations mostly for spheroids with diameter clearly smaller than before treatment ($d_{T}\leq 370-400\ \mu$m). For reference, note that the resolution of the images is $2.04\ \mu\text{m}/\text{px}$, i.e., a deviation of $20\ \mu\text{m}$ corresponds to $10$px ($<1$% deviation with respect to the original 1300 $\times$ 1030 image) or $5$px in the half-resolution image for application of the U-Net.
Only a few images with larger spheroids exhibit JCD bigger than $0.2$. Most of these are ambiguous cases, with two spheroids attached to each other, which are inconsistently recognized as either one or two spheroids even by the human, see examples with JCD$=0.28-0.34$ and $0.42$ in App. Fig. A.4. While the training data set did not contain such double-spheroid cases, the optimized U-Net often segments them correctly, see examples with JCD$=0.08,0.12$ in SI Figs. A.2 and A.3, respectively. Apart from this particular case, most images with larger deviations between the segmentations refers to smaller spheroids with surrounding cell debris. Note that smaller spheroids without cell debris, i.e., untreated or shortly after treatment, do not exhibit such deviations but are accurately segmented. To thoroughly examine these larger deviations in more detail, three biological experts independently segmented 101 randomly selected images with spheroid diameter $d_{T}\leq 400\ \mu$m. For these 101 images the JCDs between the humans and the U-Net and between different humans are compared in Fig. 5. We find that on average the discrepancies between humans and U-Net are comparable to the variations across segmentations from different humans. This implies that these images represent biologically difficult, unclear, or ambiguous cases. It suggests that the larger JCD observed for some smaller spheroids rather reflects this uncertainty and not a low performance of the U-Net. Accordingly, the evaluation in Fig. 4 refers to a worse-case scenario, as for each spheroid, the biological expert intentionally segmented all images over time, while in practice, segmentation is only required and performed for a fraction of these images, in particular for clearly distinguishable and larger spheroids.
Finally, we also test the automatic segmentation with the optimized U-Net on eight published test data sets from previous deep-learning approaches Lacetal2021 ; IvaParWalAleAshGelGar2014 ; PicPeiStePyuTumTazWevTesMarCas2023 , see A.3. Note that the total of 496 images originates from different conditions, including brightfield and fluorescence microscopy, RGB and 16-bit gray-scale, and different microscopes, magnifications, image resolutions, and cell models, but do not contain significant cell debris. While our automatic segmentation performs well on roughly half of the data sets, sometimes surpassing the original model corresponding to the data set, two types of images turn out problematic: (i) images with ambiguous ground truth and (ii) images on which the spheroids appears semi-transparent, with individual cells being visible throughout the spheroid. However, classical segmentation techniques work sufficiently well for both types of images (i) and (ii), making the use of deep-learning approaches in these cases unnecessary. In detail, images of type (i) contain spheroids with peculiar intensity profiles, i.e., with a compact sphere-like core surrounded by a flat patch. The corresponding published ground truth just defines the outer boundary of these patches as the boundary of the spheroid, while our automatic segmentation restricts the spheroid to the compact core. Either choice may be unreasonable depending on the goal of the analysis. The outer boundary of the patches allows computation of the projected area of the spheroid while ignoring the considerable variations in thickness encoded in the intensity across the spheroid. Accordingly, it is not suited to estimate the spheroid volume necessary to evaluate volume growth. Presumably, the compact core contains most of the three-dimensional organized cells and cellular volume. It is the main component exhibiting metabolic gradients affecting therapy response RifHeg2017 ; LeeGriHarMcI2016 ; HirMenDitWesMueKun2010 ; FriEbnKun2007 ; KunFreHofEbn2004 , but certainly, it underestimates the total cell volume. Segmentation of such incompletely formed spheroids is debatable, even biologically, as the formation of proliferative and metabolic gradients remains ambiguous and may depend on the question pursued with the experiment. For images of type (ii), the apparent visibility of individual cells throughout the spheroid may be due to its small size or the chosen microscopy method.
V Discussion
We develop an automatic segmentation for images of 3D tumor spheroids both with and without (radio)therapy. We systematically validate the automatic segmentation on larger, independent Hold-out test data sets of two cell types of head-and-neck cancer spheroid types, including combinations with hyperthermia treatment. For most images, we find excellent overlap between manual and automatic segmentation. These include clearly discernable spheroids and the previously neglected cases of spheroids critically obscured by cell debris. For images showing poor overlap of the segmentations, we demonstrate that this error is comparable to the variations between segmentations from different biological experts (inter-observer variability). This suggests that considerable deviations between automatic and manual segmentations do not necessarily reflect a low performance of the former but rather a general uncertainty or ambiguity in spheroid identification.
While the accuracy of (spheroid) segmentations is usually only quantified in terms of the JCD or Jaccard index (IoU), we choose to additionally report the corresponding implications on the spheroid diameter derived from the segmentation by $\Delta r$ and RDD. This makes the evaluation more intuitive for biological experts, as the spheroid diameter is the central metric for analysis. For instance, the spheroid volume required for growth curve documentation and growth delay is estimated from this diameter, which is connected to the projected area of the spheroid. We do not explicitly translate the evaluation to the corresponding spheroid area and volume. However, this may be estimated from basic scaling arguments, i.e., area and volume scale square and cubic with the diameter. The circularity of the spheroid, as quality control for its desired spherical shape, is taken into account by the RCD. All of these metrics will converge to their correct value, when the JCD goes to 0 (or Jaccard index to 1).
In fact, it should be pointed out that the imprecisions of the segmentation have a smaller effect on the quantification of therapy response than suggested by the reported JCD: As most informative parameters in long-term spheroid-based assays are the average diameters, volumes, and circularities of the spheroids. Thus, the relevant metric to estimate the error of the automatic analysis is not the JCD but the RDD and RCD, which are considerably smaller as even an imprecise segmentation can result in the correct size or shape of the spheroid. Furthermore, higher deviations are primarily observed for for treated 3D cultures with diameters below the initial diameter of standard spheroids $d_{T}=370-400\ \mu$m before treatment. These are usually images of spheroids in the final steps of detachment or before cell re-aggregation, growth recovery, and spheroid relapse. In experimental practice, such cases are typically not segmented as they are less relevant for the analytical endpoints. For instance, computation of growth delay requires accuracy of the spheroid segmentation immediately before treatment (for images without obscuring cell debris) and at large spheroid sizes $d_{T}>600\ \mu$m, while the images in between have no impact on the growth delay. Moreover, the validation in Fig. 4 is based on images from the time series of relapsed and controlled spheroids. However, in practical routine, segmentations of complete time series are only performed for growth curves and growth delay assessment in spheroid populations at 100 % growth recovery. The average JCD is substantially smaller if the validation is restricted to these cases.
It is important to note that the initial automatic segmentation is driven by images of only one HNSCC spheroid model (FaDu) with and without treatment. However, we show that it works equally well for spheroids from another cell type (SAS), although these SAS spheroids display a different peripheral shape during regrowth after treatment. Beyond the systematic validation highlighted herein, the automatic segmentation has been continuously tested and applied by several biological/biomedical experts and researchers for over a year during their ongoing experiments and to retrospectively reanalyze MCTS from earlier studies. So far, it has been reported that resegmentation is only necessary in a small fraction of cases, mainly due to optical artifacts, like out-of-focus images. Overall, the segmentation has in the meantime been successfully applied to numerous untreated multicellular spheroid types of different tumor entities and cell lines, respectively (FaDu, SAS - head and neck; Panc-02(mouse), Panc-1, PaTu 8902 - pancreas; HCT-116, HT29 - colon; A549, NCI-H23, NCI-H460 - lung; BT474 - breast; LNCap, DU145 - prostate; U87-MG, U138-MG, U251-MG - brain/GBM; Hek293 - kidney). These spheroids ranged between $200-1000\ \mu$m in diameter and images were taken at different magnifications and resolutions (e.g., $1300\times 1030$, $1388\times 1040$ $1920\times 1440$, $1920\times 1216$ and $1.6-2.6\ \mu$m/pixel) as single or Z-stack images at diverse microscopic devices (Axiovert 200M, AxioObserver Z1 - both from Zeiss; BioTek Cytation 5 Cell Imaging Multimode Reader - Agilent). Many more spheroid types are in the pipeline for implementation. The developed automatic segmentation has also already been applied to contour the images of selected spheroid types, e.g., FaDu, SAS, Panc-02 or DU145, after various treatments such as radiotherapy (X-ray, proton), hyperthermia, chemotherapy and combinatorial treatment.
We also test the automatic segmentation on eight published test data sets from previous deep-learning approaches Lacetal2021 ; IvaParWalAleAshGelGar2014 ; PicPeiStePyuTumTazWevTesMarCas2023 . We observe a high accuracy, except for special cases, which, however, turn out to be well segmentable by classical techniques. While the automatic segmentation has been continuously tested for over a year on different cell lines, microscopes, and treatments, it does not necessarily generalize to arbitrary experimental conditions and imaging, which will be the focus of future improvements. However, this problem of domain shift, which can always arise when a model is applied to data sets with a different data distribution than the training data, is also an ongoing challenge for deep-learning models focused on images with clear spheroids and clean background GarDomHerMatPas2024 . In contrast, the focus of this study is the inclusion of the case of treated spheroids surrounded by severe cell debris, which is frequent after treatment but often neglected in automatic segmentation.
Note that the automatic segmentation’s accuracy and reliability may not be exclusive to the finally chosen network architecture and hyperparameters. Indeed, we find that two very different network architectures, U-Net and HRNet, achieve very similar performance, which is consistent with previously reported optimizations of deep-learning models  (Lacetal2021, , Tab. 2) (PicPeiStePyuTumTazWevTesMarCas2023, , Tab. 3). The segmentation’s accuracy also seems relatively insensitive to the choice of several hyperparameters, e.g., loss function, optimizer function, and, to some extent, the resize factor. Instead, it is plausible, that applying challenging training data, i.e., images with extensive, severe debris, is crucial for the performance of the resulting automatic segmentation. The network trained primarily on such images also works on images with clearly visible spheroids. Hence, the data sets compiled and annotated for this work are also publicly provided to support future improvements of spheroid segmentations. This data set can facilitate the development of more individual, more customized models, e.g., using the recently introduced nnU-Net tool IseJaeKohPetMai2021 .
The automatic segmentation can be incorporated into existing or future tools for spheroid analysis or medical image analysis as basis for further feature extraction including perimeter, complexity, and multiparametric analysis ZhuCheWanTanHeQinYanJinYuJinLiKet2022 . The segmentation provides a basis for the development of an automatic classification of spheroid image time series into control and relapse and could support machine learning methods to forecast tumor spheroid fate early. To make the deep-learning model available, we provide a minimal tool with graphical user interface. The ONNX Runtime Ord2021 is used to compile the model and to take it into production.
Availability and requirements
Segmentation tool is published as open-source repository https://igit.informatik.htw-dresden.de/aagef650/spheroidsegdedeb (additionally download the model from https://wwwpub.zih.tu-dresden.de/~s6079592/SEGMODEL.onnx and copy it into the model/ folder) or latest version for download https://wwwpub.zih.tu-dresden.de/~s6079592/. Requirements and usage are documented in the provided README and some example images for testing are included in the corresponding folder. Training, validation, and test data used for this work will be available after peer-reviewed publication as well as the code to train and evaluate the model (see above repo training_scripts/). All networks optimized for different hyperparameters are available in the .pth format at https://wwwpub.zih.tu-dresden.de/~s6079592/models/.
Acknowledgement
We are grateful to M. Wondrak, S. Al-Jamei, N. El-Refai, R. Joseph, and S. L. Prieto for sharing their expertise on tumor spheroid contouring.
Funding
The authors acknowledge that this research has been co-financed by the EU, the European Social Fund (ESF), and by tax funds on the basis of the budget passed by the Saxon state parliament (project SAB-Nr. 100382145) and the Bundesministerium für Bildung und Forschung (BMBF 16dkwn001a/b). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Declaration of Competing Interest
The authors declare that they have no competing interests.
Appendix A Supplemental figures
(a) Untreated spheroid w/o cell debris
(b) Treated spheroid with cell debris
Figure A.1: Example images illustrating the current challenge of segmenting tumor spheroids: (a) While previously developed deep-learning models generalize excellently to new image data with well-distinguished, unobscured spheroids, typical for untreated cultures, (b) these models fail for cases of detached/relapsing spheroids after radiotherapy (one of the most common cancer treatments) due to debris and dead cells obscuring the spheroid. Shown are representative images of FaDu spheroids (a) without treatment and (b) several days after radiotherapy. Different segmentations are indicated by their outer contours. In particular, the manually set ground truth (green solid line) is compared to the four most recent deep-learning models provided by SpheroidJ Lacetal2021 (HRNet-Seg with HRNet W30 backbone, yellow close-dashed line), SpheroScan AksKatAbeSheBesBurBigAdaMonGhe2023 (Region-based Convolutional Neural Network trained on images from IncuCyte Live-Cell Analysis System (purple dashed line) and ordinary microscope (torquoise dashed line) with recommended threshold 0.8), and AnaSP PicPeiStePyuTumTazWevTesMarCas2023 (ResNet18, red dotted line). Only the U-Net presented in this manuscript (blue solid line) performs sufficiently in both scenarios. Size of each image corresponds to $2650\ \mu\text{m}\times 2100\ \mu\text{m}$ with a resolution of 1300 $\times$ 1030 pixel.
Table A.2: Statistical results reflecting the current challenge of segmenting tumor spheroids, corroborating App. Fig. A.1: While previously developed deep-learning models generalize excellently to new image data with well-visible, unobscured spheroids (top line), these models fail for typical cases of detached/relapsing spheroids with debris of dead cells (bottom lines). Intersection over Union (IoU or Jaccard index) is reported (average (median) $\pm$ standard deviation) for each of the four most recent deep-learning models (provided by SpheroidJ Lacetal2021 , SpheroScan AksKatAbeSheBesBurBigAdaMonGhe2023 (with recommended threshold 0.8), and AnaSP PicPeiStePyuTumTazWevTesMarCas2023 ) and each of our data sets. Only the U-Net presented in this manuscript performs sufficiently in both scenarios. Bold values highlight the optimum for each dataset.
Data set #Images Intersection over Union (IoU = 1-JCD)
U-Net SpheroidJ Lacetal2021 SpheroScan (Incu) AksKatAbeSheBesBurBigAdaMonGhe2023 SpheroScan (Micro) AksKatAbeSheBesBurBigAdaMonGhe2023 AnaSP PicPeiStePyuTumTazWevTesMarCas2023
Spheroids without
cell debris
200 $\mathbf{0.97(0.97)\pm 0.01}$ $0.89(0.90)\pm 0.05$ $0.90(0.94)\pm 0.16$ $0.72(0.92)\pm 0.38$ $0.81(0.87)\pm 0.22$
Training 883 $\mathbf{0.94(0.95)\pm 0.05}$ $0.51(0.53)\pm 0.27$ $0.68(0.86)\pm 0.34$ $0.48(0.59)\pm 0.37$ $0.46(0.44)\pm 0.28$
Validation 108 $\mathbf{0.95(0.96)\pm 0.04}$ $0.31(0.29)\pm 0.19$ $0.59(0.80)\pm 0.39$ $0.36(0.26)\pm 0.32$ $0.28(0.25)\pm 0.19$
Test 104 $\mathbf{0.94(0.96)\pm 0.06}$ $0.41(0.38)\pm 0.27$ $0.66(0.88)\pm 0.36$ $0.48(0.50)\pm 0.38$ $0.35(0.33)\pm 0.24$
Table A.3: Further validation of the trained U-Net on wide variety of published test data sets from previous deep-learning models. While the trained U-Net performs well on on roughly half of the data sets or 38% of the images (with average IoU above 0.8), sometimes surpassing the original model corresponding to the data set, two types of images turn out problematic: (i) images with ambiguous ground truth (20% of images) for which the U-Net may actually segment reasonably, see main text for detailed discussion, and (ii) images on which the spheroids appears semi-transparent (42% of images), with individual cells being visible throughout the spheroid, potentially due to its small size or the chosen microscopy method. However, classical segmentation techniques work sufficiently well for both types of images (i) and (ii), making the use of deep-learning approaches in these cases unnecessary. To demonstrate this, the last two columns report the performance of the classical approach from the original publication and simple Otsu thresholding (sometimes after some image erosion for (ii)) performed by us. Note that the notation for data sets from Refs. IvaParWalAleAshGelGar2014 ; Lacetal2021 stands for brightfield/fluorescence microscopy (B/F), Nikon Eclipse/Leica DMi8/Olympus microscope (N,L,O), 2x/5x/10x magnification (2/5/10), and suspension/collagen culture (S/C). Bold values highlight the optimum for each dataset.
Data set #Images Comments Intersection over Union (IoU = 1-JCD)
deep-learning model classical segmentation
Original U-Net Original Otsu thresholding
BO10S IvaParWalAleAshGelGar2014 66 $0.92\pm 0.03$ Lacetal2021 $\mathbf{0.97(0.98)\pm 0.05}$ $0.94\pm 0.03$ Lacetal2021
BN10S Lacetal2021 21 $\mathbf{0.97\pm 0.01}$ $0.89(0.89)\pm 0.01$ $0.95\pm 0.01$ $0.84(0.85)\pm 0.13$
84 ambiguous ground truth $0.61(0.62)\pm 0.15$
BN2S Lacetal2021 154 semi-transparent $\mathbf{0.96\pm 0.01}$ $0.30(0.27)\pm 0.22$ $0.94\pm 0.02$ $0.87(0.95)\pm 0.27$
BL5S Lacetal2021 50 semi-transparent $0.75\pm 0.25$ $0.37(0.13)\pm 0.39$ $0.64\pm 0.30$ $\mathbf{0.85(0.93)\pm 0.20}$
FN2S Lacetal2021 30 $0.78\pm 0.20$ $0.91(0.94)\pm 0.08$ $\mathbf{0.82\pm 0.17}$
4 semi-transparent 00 $0.83(0.83)\pm 0.02$
FL5CLacetal2021 4 $0.71\pm 0.30$ $0.92(0.91)\pm 0.01$ $0.67\pm 0.17$
15 ambiguous ground truth $0.45(0.45)\pm 0.16$ $\mathbf{0.80(0.85)\pm 0.11}$
FL5S Lacetal2021 50 $0.70\pm 0.26$ $\mathbf{0.91(0.91)\pm 0.05}$ $0.89\pm 0.07$
AnaSP PicPeiStePyuTumTazWevTesMarCas2023 16 $\sim 0.92$ $0.97(0.98)\pm 0.02$
2 semi-transparent $0.01(0.01)\pm 0.01$ $0.97(0.97)\pm 0.00$
Figure A.2: Selection of representative images as optical reference for JCD $\leq 0.1$ with automatic segmentation from the optical U-Net (orange) and manual segmentation from biological expert H2 (green). From the tested images, $52\%$ fall into this range of JCD $\leq 0.1$ and $67\%$ when only spheroids beyond the standard size $d_{T}>400\ \mu$m are considered. Size of each image corresponds to $2650\ \mu\text{m}\times 2100\ \mu\text{m}$. Note that the U-Net is trained on an independent manual segmentations from another biological expert H1. Images are selected from the extended validation data set (Fig. 4), in particular, from the subset of larger, better discriminated spheroids at intermediate levels of debris for illustration. Note that in the example for JCD$=0.08$, the two attached spheroids are correctly segmented, although the training data set does not contain such cases.
[…]
Figure A.4: Selection of representative images as optical reference for larger deviations JCD$>0.2$ of the automatic segmentation from the optical U-Net (orange) and manual segmentation from biological expert H2 (green), analogous to App. Fig. A.2. From the tested images, $26\%$ fall into the range JCD$>0.2$, but merely $10\%$ when only spheroids beyond the standard size $d_{T}>400\ \mu$m are considered. Note that deviations at larger spheroids are often due to cases of double-spheroids, e.g., JCD$=0.27,0.28,0.29,0.34,0.42$, which are inconsistently recognized as either one or two spheroids even by the human.
Backbone JCD RDD RCD ISF ASF
ResNet 18 0.192 0.131 0.142 0.074 0.028
$\pm$ 0.282 0.275 0.258 0.025 0.016
ResNet 34 0.100 0.046 0.055 0.009 0.000
$\pm$ 0.133 0.116 0.100 0.009 0.000
xResNet 18 0.312 0.193 0.376 0.074 0.880
$\pm$ 0.303 0.300 0.268 0.025 0.031
xResNet 34 0.228 0.155 0.174 0.083 0.065
$\pm$ 0.297 0.292 0.268 0.027 0.024
SqueezeNet1.0 0.317 0.215 0.316 0.130 0.491
$\pm$ 0.292 0.315 0.282 0.032 0.048
SqueezeNet1.1 0.237 0.148 0.207 0.056 0.352
$\pm$ 0.281 0.271 0.237 0.022 0.046
VGG 11-bn 0.170 0.096 0.247 0.028 0.315
$\pm$ 0.210 0.193 0.186 0.016 0.045
VGG 13-bn 0.161 0.089 0.248 0.019 0.315
$\pm$ 0.193 0.170 0.173 0.013 0.045
VGG 16-bn 0.157 0.088 0.256 0.028 0.287
$\pm$ 0.185 0.173 0.162 0.016 0.044
VGG 19-bn 0.118 0.057 0.213 0.000 0.287
$\pm$ 0.117 0.090 0.133 0.000 0.044
AlexNet 0.375 0.280 0.361 0.074 0.398
$\pm$ 0.265 0.346 0.219 0.025 0.047
Figure A.5: Values of evaluation metrics for different backbones of the U-Net. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Optimizer: Adam, Loss: Cross-Entropy, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: No. Overall the ResNet 34 achieved the best results and is picked as the backbone for the U-Net.
Backbone JCD RDD RCD ISF ASF
W18 0.290 0.185 0.196 0.065 0.241
$\pm$ 0.311 0.291 0.249 0.024 0.041
W30 0.211 0.139 0.130 0.056 0.028
$\pm$ 0.271 0.266 0.231 0.022 0.016
W32 0.289 0.201 0.258 0.102 0.111
$\pm$ 0.320 0.330 0.290 0.029 0.030
W48 0.165 0.064 0.131 0.009 0.167
$\pm$ 0.132 0.111 0.143 0.009 0.036
Figure A.6: Values of evaluation metrics for different backbones of the HRNet. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Optimizer: Adam, Loss: Cross-Entropy, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: No. Overall the W48 achieved the best results and is picked as the backbone for the HRNet.
Extension JCD RDD RCD ISF ASF
None 0.151 0.058 0.069 0.000 0.028
$\pm$ 0.170 0.086 0.084 0.000 0.016
Transfer learning 0.078 0.028 0.043 0.000 0.000
$\pm$ 0.058 0.034 0.037 0.000 0.000
Data augmentation 0.121 0.048 0.050 0.000 0.000
$\pm$ 0.130 0.072 0.066 0.000 0.000
Transfer learning & 0.074 0.026 0.041 0.000 0.000
Data augmentation $\pm$ 0.053 0.027 0.035 0.000 0.000
Figure A.7: Values of evaluation metrics for different extensions of the training data set, which is used by the U-Net. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: ResNet 34, Optimizer: Adam, Loss: Cross-Entropy, Resize factor: 1/2. The accuracy of the U-Net is highest when transfer learning and data augmentation is used.
Extension JCD RDD RCD ISF ASF
None 0.169 0.078 0.071 0.000 0.019
$\pm$ 0.196 0.135 0.103 0.000 0.013
Transfer learning 0.109 0.041 0.062 0.000 0.009
$\pm$ 0.123 0.069 0.082 0.000 0.009
Data augmentation 0.113 0.040 0.043 0.000 0.000
$\pm$ 0.115 0.052 0.040 0.000 0.000
Transfer learning & 0.067 0.022 0.043 0.000 0.000
Data augmentation $\pm$ 0.046 0.022 0.032 0.000 0.000
Figure A.8: Values of evaluation metrics for different extensions of the training data set, which is used by the HRNet. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: W48, Optimizer: Adam, Loss: Cross-Entropy, Resize factor: 1/2. The accuracy of the HRNet is highest when transfer learning and data augmentation is used.
Resize factor JCD RDD RCD ISF ASF
1/4 0.085 0.029 0.121 0.000 0.000
$\pm$ 0.071 0.042 0.101 0.000 0.000
1/2 0.074 0.026 0.041 0.000 0.000
$\pm$ 0.053 0.027 0.035 0.000 0.000
3/4 0.076 0.028 0.046 0.000 0.000
$\pm$ 0.075 0.047 0.036 0.000 0.000
1 0.163 0.094 0.140 0.028 0.130
$\pm$ 0.232 0.231 0.192 0.016 0.032
Figure A.9: Values of evaluation metrics for different image sizes. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: ResNet 34, Optimizer: Adam, Loss: Cross-Entropy, Transfer learning: Yes, Data augmentation: Yes. The U-Net can achieve the highest accuracy when the original image size is reduced by half.
Resize factor JCD RDD RCD ISF ASF
1/4 0.091 0.035 0.125 0.000 0.009
$\pm$ 0.090 0.058 0.106 0.000 0.009
1/2 0.067 0.022 0.043 0.000 0.000
$\pm$ 0.046 0.022 0.032 0.000 0.000
3/4 0.094 0.037 0.045 0.000 0.019
$\pm$ 0.118 0.086 0.045 0.000 0.013
1 0.264 0.171 0.226 0.065 0.269
$\pm$ 0.282 0.274 0.235 0.024 0.043
Figure A.10: Values of evaluation metrics for different image sizes. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: W48, Optimizer: Adam, Loss: Cross-Entropy, Transfer learning: Yes, Data augmentation: Yes. The HRNet can achieve the highest accuracy when the original image size is reduced by half.
Loss function JCD RDD RCD ISF ASF
Cross-Entropy 0.074 0.026 0.041 0.000 0.000
$\pm$ 0.053 0.027 0.035 0.000 0.000
Focal 0.076 0.025 0.058 0.000 0.000
$\pm$ 0.059 0.031 0.042 0.000 0.000
Dice 0.067 0.022 0.041 0.000 0.000
$\pm$ 0.048 0.023 0.035 0.000 0.000
Figure A.11: Values of evaluation metrics for different loss functions, used by the U-Net. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: ResNet 34, Optimizer: Adam, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes. The U-Net can achieve the highest accuracy when the Dice loss is used.
Loss function JCD RDD RCD ISF ASF
Cross-Entropy 0.067 0.022 0.043 0.000 0.000
$\pm$ 0.046 0.022 0.032 0.000 0.000
Focal 0.067 0.022 0.040 0.000 0.000
$\pm$ 0.047 0.023 0.032 0.000 0.000
Dice 0.100 0.038 0.053 0.009 0.000
$\pm$ 0.125 0.099 0.105 0.009 0.000
Figure A.12: Values of evaluation metrics for different loss functions, used by the HRNet. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: W48, Optimizer: Adam, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes. The HRNet can achieve the highest accuracy, if the Cross-Entropy or the Focal loss is used.
Optimizer function JCD RDD RCD ISF ASF
Adam 0.067 0.022 0.041 0.000 0.000
$\pm$ 0.048 0.023 0.035 0.000 0.000
RAdam & Lookahead 0.066 0.024 0.038 0.000 0.000
$\pm$ 0.046 0.026 0.033 0.000 0.000
Figure A.13: Values of evaluation metrics for different optimizers, used by the U-Net. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: ResNet 34, Loss: Dice, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes. The U-Net can achieve the highest accuracy when the optimization is done by RAdam combined with Lookahead.
Optimizer function JCD RDD RCD ISF ASF
Adam 0.067 0.022 0.043 0.000 0.000
$\pm$ 0.046 0.022 0.032 0.000 0.000
RAdam & Lookahead 0.106 0.047 0.061 0.000 0.000
$\pm$ 0.142 0.112 0.083 0.000 0.000
Figure A.14: Values of evaluation metrics for different optimizers, used by the HRNet. Bold values highlight the optimum in each column. The other hyperparameters are fixed at Backbone: W48, Loss: Cross-Entropy, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes. The HRNet can achieve the highest accuracy when the optimization is done by Adam.
Model JCD RDD RCD ISF ASF
U-Net 0.062 0.020 0.040 0.000 0.010
$\pm$ 0.060 0.026 0.034 0.000 0.010
HRNet 0.072 0.028 0.036 0.000 0.010
$\pm$ 0.100 0.080 0.029 0.000 0.010
Figure A.15: Evaluation of the segmentation with the optimized U-Net and HRNet models on the test data set shows higher accuracy of the U-Net. The optimal hyperparameter configuration for the U-Net is Backbone: ResNet 34, Optimizer: RAdam & Lookahead, Loss: Dice, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes. The final hyperparameter configuration for the HRNet is Backbone: W48, Optimizer: Adam, Loss: Cross-Entropy, Resize factor: 1/2, Transfer learning: Yes, Data augmentation: Yes.
(a) Image with segmentations
(b) Segmentations with mismatched (gray) areas
(c) Corresponding circle with rim of width $\Delta r$
Figure A.16: Illustration of the average radial error $\Delta r$ defined in Eq. (3) (example with large JCD$>0.2$ for visibility). (a) Zoom of image from App. Fig. A.4 with JCD$=0.24$ with corresponding automatic (orange line containing predicted area $P$) and manual segmentation (green line containing target area $T$). (b) Segmentations from (a) with mismatched areas shown in gray (missing area $T\setminus P$ in bright and dark gray, additional area $P\setminus T$ in intermediate gray). (c) Circle (green line) and added rim (gray areas) with areas corresponding to (b): Area within green circle is equal to area within manual segmentation in (b) and gray areas within the adjacent rim are equal to corresponding mismatched areas in (b). Then the radial thickness of the rim is equal to the average radial error $\Delta r$ defined in Eq. (3).
(a) RDD over spheroid diameter
(b) Comparison diameters from manual and automatic segmentation
Figure A.17: Validation based on diameter of automatic segmentation with the optimized U-Net on larger, independent data sets, analogous to Fig. 4. (a) RDD shows that majority of deviations are below $10\%$ and higher deviations occur mostly below $d_{T}<400\ \mu$m. (b) Direct comparison of diameters (gray points) resulting from automatic segmentation (predicted diameter) and manual segmentation (target diameter) shows high accuracy (red line represents perfect match) with larger deviations mostly at $d_{T}<400\ \mu$m.